OpenStudy (anonymous):

Solve 9k^2-4k=0 pls help

OpenStudy (saifoo.khan):

9k^2-4k=0 take commons, k(9k -4) = 0 k = 0, 9k - 4 = 0 k = 4/9

OpenStudy (anonymous):

thank you so much